• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 16:15:25 on July 5,2010

    Entry number 330251

    keyword=Day Shift Summary

    Shift summary
    -----------------

    8.00 : Beam was down at the start of the shift.
    9.20 : Still waiting for the beam...
    9.30 : Controlled access to adjust delays in Cerenkov PMTs(Bogdan, Eric)
    10.20 : Finished controlled access. Still no beam.
    10.40 : MCC informs that beam is ready to be delivered. Asked for 5uA.
    11.00 : Started taking data with 5uA on Pb.
    11.24 : Q1 tripped. Reset Q1
    12.20 : Finished controlled access. Started data-taking.
    12.40 : Target moved to Ta
    4.00 : move to Pb 1 target

    Run summary:

    Run Target Current comments
    -------------------------------------
    1729 Pb 5uA
    1730 Pb 10uA Q1 tripped. stopped run
    1731 Pb 10uA Deadtime too high
    1732 Pb 10uA 1M events
    1733 Ta 10uA T1 triggers
    1734 Ta 1uA T1 triggers, raster off
    1735 Ta 1uA T3 triggers, raster off
    1736 Ta 25uA T6 triggers
    1737 Ta - junk(test)
    1738 Ta 25uA T6 triggers
    1739 Ta 10uA T3 triggers
    1740 Ta 10uA T4 triggers (beam tripped)
    1741 Ta 10uA T4 continues
    1742 Ta 2uA T4 triggers
    1743 Ta 2uA T3
    1744 Ta 2uA T3
    1745 Ta 2uA T4
    1746 Ta 2uA T4
    1747 Ta 2uA T3
    1748 Ta 2uA T3